# PRoTECT: Parallelized ConstRuction of SafeTy BarriEr Certificates for Nonlinear Polynomial SysTems

### v1

@@ 8. Conclusion
This work introduced PRoTECT, a pioneer software tool utilizing SOS optimization to explore polynomial-type BCs for verifying safety properties across four classes of dynamical systems: dt-SS, dt-DS, ct-SS, and ct-DS. The tool is developed in Python and incorporates a user-friendly GUI to enhance its usability. Additionally, PRoTECT offers parallelization to concurrently search for BCs of different degrees, ensuring an efficient construction. In the future, PRoTECT will be expanded to incorporate reachability and reach-while-avoid specifications, along with designing controllers using SOS optimization techniques.
@@ A.1. Discrete-time Stochastic Systems (dt-SS)
$\Sigma_{d}^{\varsigma}\!:x(k+1)=(1-\beta-\theta\nu)x(k)+\theta T_{h}\nu(k)+\beta T_{e}+R\varsigma(k),$
@@
A.1.2. 2D Van der Pol oscillator.
Figure 3. $2$D Van der Pol oscillator system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(0,4)$. The uniform noise has bound [-0.1, 0.1] in both dimensions.
A.1.3. Two-Tank System.
Figure 4. $2$D Two-Tank system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(2.2,2.2)$.
@@ A.2. Discrete-time Deterministic Systems (dt-DS)
Figure 5. $2$D Room system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(19.5,19.5)$.
@@
A.2.3. 2D DC Motor.
Figure 6. $2$D DC Motor system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(0.3,0.2)$.
We consider the following DC motor case study [3]
@@ A.3. Continuous-time Stochastic Systems (ct-SS)
Figure 7. $2$D linear system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(-0.25,2.75)$.
@@
A.3.3. 2D Nonlinear System.
Figure 8. $2$D nonlinear system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(0,0)$.
We consider the following nonlinear system [25]
$\Sigma_{c}^{\delta}:\begin{cases}\mathbf{d}{x}_{1}(t)=x_{2}(t)\mathbf{d}t,\\ \mathbf{d}{x}_{2}(t)=(-x_{1}(t)-x_{2}(t)-0.5x_{1}^{3}(t))\mathbf{d}t+\delta\mathbf{d}\mathbb{W}_{t},\end{cases}$
A.3.4. Benchmark - High Order 4.
@@ A.4. Continuous-time Deterministic Systems (ct-DS)
A.4.1. 2D Jet Engine.
Figure 9. $2$D Jet Engine system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black starting from an initial state $x_{0}=(0.2,0.2)$.
We consider the nonlinear Moore-Greitzer jet engine model [14]:
A.4.2. FOSSIL Benchmark - High Order 4.
@@
A.4.3. FOSSIL benchmark - High Order 6.
@@
A.4.4. FOSSIL benchmark - High Order 8.
@@
Table 2. Efficiency evaluation in BC construction for stochastic systems via PRoTECT. We compare case studies across three barrier degrees (2, 4, and 6), returning the barrier with the highest confidence. Additionally, we evaluate an optimized version using different fixed $\lambda$ values. VDP denotes the stochastic Van der Pol oscillator. All experiments were conducted on a desktop computer (Intel i9-12900).
Standard Optimized
$n$ system $\mathcal{T}$ b_degree $\gamma$ $\lambda$ $c$ $\phi$ time (sec) b_degree $\gamma$ $\lambda$ $c$ $\phi$ time (sec)
RoomTemp [23] 1 ct-SS 5 6 0.74 3.55 0.41 0.23 0.33 6 $1.1e^{-6}$ 10 $2.3e^{-7}$ 0.99 0.33
RoomTemp [23] 1 dt-SS 5 6 1.12 6.16 0.72 0.23 0.52 6 $4.4e^{-7}$ 10 $1.0e^{-7}$ 0.99 0.53
VDP [2] 2 dt-SS 5 6 327 610 90.5 0.02 14.1 6 97.5 1000 3.53 0.88 14.3
ex_lin1 [25] 2 ct-SS 5 6 5.81 11.2 0.99 0.03 1.81 6 0.34 10 0.04 0.95 1.73
ex_nonlin1 [25] 2 ct-SS 5 6 12.4 42.8 5.73 0.04 1.63 6 1.84 10 0.2 0.72 1.81
TwoTanks [28] 2 dt-SS 5 4 6.23 265 8.06 0.82 5.13 4 $1.0e^{-6}$ 10 $3.4e^{-8}$ 0.99 5.14
RoomTemp [31] 3 dt-SS 3 6 2.55 6248 1.74 0.99 1509 4 $1.1e^{-8}$ 10 $7.5e^{-9}$ 0.99 1501
hi-ord4 [1] 4 ct-SS 3 2 5.12 30.4 7.78 0.05 1309 6 $1.8e^{-3}$ 10 $1.2e^{-3}$ 0.99 1308

### v2

@@ 8. Conclusion
This work introduced PRoTECT, a pioneer software tool utilizing SOS optimization to explore polynomial-type BCs for verifying safety properties across four classes of dynamical systems: dt-SS, dt-DS, ct-SS, and ct-DS. In particular, PRoTECT is the first software tool that offer stochastic barrier certificates. The tool is developed in Python and incorporates a user-friendly GUI to enhance its usability. Additionally, PRoTECT offers parallelization to concurrently search for BCs of different degrees, ensuring an efficient construction. In the future, PRoTECT will be expanded to incorporate reachability and reach-while-avoid specifications, along with designing controllers using SOS optimization techniques.
@@ A.1. Discrete-time Stochastic Systems (dt-SS)
$\Sigma_{d}^{\varsigma}\!:x(k+1)=(1-\beta-\theta\nu)x(k)+\theta T_{h}\nu+\beta T_{e}+R\varsigma,$
@@
A.1.2. Two-Tank System.
Figure 7. $2$D Two-Tank system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(2.2,2.2)$ for $\mathcal{T}=20$, with time increments of $0.01$.
@@ A.2. Discrete-time Deterministic Systems (dt-DS)
Figure 8. $2$D Room system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(19.5,19.5)$ for a time horizon of 1000 steps.
@@ A.3. Continuous-time Stochastic Systems (ct-SS)
Figure 9. $2$D linear system, with initial and unsafe regions $X_{\mathcal{I}},X_{\mathcal{U}}$ marked by blue and red boxes, respectively. Level sets of the barrier certificate are displayed in solid blue and red dashed for $\mathcal{B}(x)=\gamma$ and $\mathcal{B}(x)=\lambda$, respectively. A trajectory of the system is shown in black with initial point $x_{0}=(-0.25,2.75)$ for a time horizon $\mathcal{T}=5$, and time increments $0.001$.
@@
A.3.3. Benchmark - High Order 4.
@@ A.4. Continuous-time Deterministic Systems (ct-DS)
A.4.1. FOSSIL Benchmark - High Order 4 (hi-ord4).
@@
A.4.2. FOSSIL benchmark - High Order 6 (hi-ord6-1).
@@
A.4.3. High Order 6 (hi-ord6-2)
We consider the following $6$-dimensional example
$\Sigma_{c}:\begin{cases}\dot{x}_{1}(t)=x_{2}(t)-100x_{3}(t),\\ \dot{x}_{2}(t)=x_{3}(t),\\ \dot{x}_{3}(t)=x_{4}(t)-100x_{5}(t),\\ \dot{x}_{4}(t)=x_{5}(t),\\ \dot{x}_{5}(t)=x_{6}(t)-100x_{1}(t),\\ \dot{x}_{6}(t)=-800x_{6}(t)-2273x_{5}(t)-3980x_{4}(t)-4180x_{3}(t)-2400x_{2}(t)-576x_{1}(t),\end{cases}$
A.4.4. FOSSIL benchmark - High Order 8 (hi-ord8-1).
@@ A.4.3. High Order 6 (hi-ord6-2)
A.4.5. High Order 8 (hi-ord8-2).
We consider the following $8$-dimensional example
$\Sigma_{c}:\begin{cases}\dot{x}_{1}(t)=&x_{2}(t)-50x_{3}(t),\\ \dot{x}_{2}(t)=&x_{3}(t)-50x_{4}(t),\\ \dot{x}_{3}(t)=&x_{4}(t)-50x_{5}(t),\\ \dot{x}_{4}(t)=&x_{5}(t)-50x_{6}(t),\\ \dot{x}_{5}(t)=&x_{6}(t)-50x_{7}(t),\\ \dot{x}_{6}(t)=&x_{7}(t)-50x_{8}(t),\\ \dot{x}_{7}(t)=&x_{8}(t)-50x_{1}(t),\\ \dot{x}_{8}(t)=&-20x_{8}(t)-170x_{7}(t)-800x_{6}(t)-2273x_{5}(t)-3980x_{4}(t)-4180x_{3}(t)\\ &-2400x_{2}(t)-576x_{1}(t),\end{cases}$
with $X=[-2.2,2.2]^{8}$, $X_{\mathcal{I}}=[0.9,1.1]^{8}$, and $X_{\mathcal{U}}=[-2.2,-1.8]^{8}$.